# Training-Free Safe Denoisers for Safe Use of Diffusion Models

### v2

@@ 5.1 Text-to-Image Generation
Table 3: Experiments on ImageNet for the specific class (Chihuahua) negation task. Top-1 is the classification accuracy of the generated samples on 999 classes, and Top-1* indicates the accuarcy on the specific class.
Method Prec $\uparrow$ Rec $\uparrow$ Top-1 $\uparrow$ Top-1 $\downarrow$
Baseline (B) 0.72 0.63 0.76 0.68
B + SR 0.59 0.54 0.01 0.0
B + Ours 0.62 0.58 0.14 0.0
Furthermore, our method significantly improves safety performance while maintaining image quality. Notably, the extent of improvement varies considerably depending on the characteristics of the prompts. For instance, with MMA-Diffusion prompts the performance of text-based baselines (like SLD) is markedly inferior compared to their performance on other prompt datasets such as Ring-A-Bell or UnlearnDiff. This discrepancy arises because MMA-Diffusion prompts lack explicit nudity information due to being part of a white-box adversarial attack, making it challenging for text-based methods to erase such content. In contrast, our approach employs purely image-based guidance, which, when combined with existing text-based methods, results in substantial performance gains from 0.88 to 0.48 in ASR on MMA-Diffusion that do not explicitly include unsafe text. Additionally, our method significantly improves the performance across all other prompt datasets, not limited to MMA-Diffusion.
Qualitative Results
Here, we present three ablation studies to evaluate the robustness and effectiveness of our method. First, Figure 4-(a) shows the effect of the number of unsafe data points on model performance. We observe that increasing the number of unsafe data points leads to better performance.
We then explore the influence of the threshold parameter $\beta_{t}$, which governs the application of the safe denoiser. For simplicity, we fixed $\beta_{t}$ across all time steps. Figure 4-(b) shows the performance exhibits a U-shaped relationship to $\beta_{t}$. Specifically, when $\beta_{t}=0$, the safe denoiser is applied to all samples $\mathbf{x}_{t}$ regardless of their safety status. Conversely, when $\beta_{t}=\infty$, the safe denoiser is not applied. At intermediate values of $\beta_{t}$, the safe denoiser is applied selectively to a certain proportion of unsafe samples $\mathbf{x}_{t}$. The U-shaped trend indicates that selectively applying the safe denoiser to unsafe samples based on an appropriate $\beta_{t}$ value is optimal, thereby balancing denoising efficacy and computational efficiency.
Finally, we assess the performance with varying the negative prompt weight in text-based methods. To establish that our approach consistently enhances performance across diverse setups when integrated with existing methodologies, we conduct a series of experiments. Figure 4-(c) shows as the weight of SLD increases from Weak to Max, SLD performs inadequately in MMA-Diffusion prompts. In contrast, our safe denoiser improves performance and widens the performance gap, underscoring its robustness and effectiveness as a superior enhancement.
5.2 Unconditional and Class Conditional Generation
In this section, we use our safe denoiser in the DMs without text inputs. Specifically, we employ experiments on FFHQ (Karras, 2019) and ImageNet (Russakovsky et al., 2015) in the $256\times 256$ resolution. We utilize the pretrained diffusion models from Chung et al. (2022) for FFHQ and Dhariwal & Nichol (2021) for ImageNet. For the experiments, we use a DPM solver (Lu et al., 2022) with 100 steps.
In FFHQ, we aim to prevent the generation of a specific sex. However, since the whole data points are used in training, we select 1K female images from CelebA (Liu et al., 2015) validation split to serve as unseen negative data, thereby establishing the negative dataset $\{\mathbf{x}^{(1)},...,\mathbf{x}^{(1000)}\}$. We then employ our safe denoiser to sample 1K images. As shown in Table 3, classification accuracy for these generated samples reveal that our method more effectively avoids the female class compared to the baseline pretrained model. According to Table 3, our algorithm generates more male images compared to Sparse Repellency (SR) while achieving a lower FID score. This suggests that the images produced by SR are of lower quality compared to those generated by ours, leading to increased confusion for the classifier.
In ImageNet, we focus on negating a specific Chihuahua class during generation. We select the validation set of Chihuahua class as the negative images. We generate 50 samples per class and classify samples from 999 classes by a classifier (Dhariwal & Nichol, 2021) and report the accuracy by Top-1. Also, we measure the Top-1 accuracy of 50 samples from Chihuahua class, reporting it by Top-1* in Table  3. From the result, we note that our method excels generating other 999 classes, while SR cannot generate images from those 999 classes. To evaluate the overall quality, Table  3 further report the precision (sample accuracy) and recall (sample diversity) (Kynkäänniemi et al., 2019) over 50K samples, indicating that our method is better than SR in negating a specific class.
6 Conclusion
@@
We have addressed significant safety challenges in DMs, particularly concerning the generation of NSFW content and the inadvertent reproduction of sensitive data. We introduce the safe denoiser, a novel approach that modifies the sampling trajectories of DMs to adhere to theoretically safe distributions, thereby ensuring the generation of appropriate and authorized content. Experimental results demonstrate that the safe denoiser achieves state-of-the-art performance in tasks such as concept erasing, class removal, and unconditional image generation. Ultimately, this work provides a robust and scalable solution for mitigating safety risks in generative AI, paving a way for their responsible and ethical applications.
Impact Statements
@@ Appendix A Proof
Appendix B Experimental Details and Additional Results
B.1 Implementation Details
Text-to-Image Generation
As outlined in the manuscript, we conduct the Text-to-Image experiment using SDv1.4, following the same model as the baselines for generating images from text, as referenced in (Schramowski et al., 2023; Wu et al., 2024; Gong et al., 2024; Yoon et al., 2024). To ensure consistency, we adopt the generation procedure described in each baseline. Preliminary observing the sensitivity of nudity-related content, we employ the DDPM scheduler (Ho et al., 2020). For a fair comparison, we maintain the same number of inference steps, specifically $50$, aligning with the official implementations of both SLD and SAFREE, which also use 50 inference steps.
@@ B.1 Implementation Details
$K(x,x^{\prime})=\exp\left(-\frac{\lVert x-x^{\prime}\rVert^{2}}{2\sigma^{2}}\right)$ (B.1)
For the bandwidth parameter $\sigma$, we set a value of 1.0 for SLD and 3.15 for SAFREE. Additionally, in case of SAFREE, we apply a scaling factor $\eta=0.33$, whereas for SLD, we use $\eta=0.03$ to regulate the strength of the repellency in Eq. (3.3). For reference images, we utilize a total of 515 images sourced from the I2P dataset (Schramowski et al., 2023), which were generated using SDv1.4. As stated in the manuscript, these reference images meet the criterion of having a nude class probability above 0.6, as determined by Nudenet. Sample images are shown below.
Figure B.1: Samples of reference images by I2P dataset
@@
Empirically, we introduce a heuristic in which the proposed Safe Denoiser is applied within critical timesteps $C=[780,...,1000]$. In the early stages of diffusion, denoising process primarily establishes global structures rather than intricate details, while the later stages focus on refining fine-grained features. Since our approach aims to prevent the generation of globally harmful images rather than enhancing image quality or detail, we apply the denoiser at these later timesteps.
@@
Next, we briefly introduce the baseline models used in our experiments. The first two approaches serve as comparisons for unlearning-based safe diffusion models (Gandikota et al., 2023a; Gong et al., 2024). Specifically, we evaluate Erased Stable Diffusion (ESD) (Gandikota et al., 2023a) as a representative method. More recently, reliably trained safe diffusion (RECE) models have demonstrated improved performance, particularly in reducing the attack success rate (Gong et al., 2024). In addition to these unlearning-based approaches, we also include SLD and SAFREE as training-free safe diffusion models (Schramowski et al., 2023; Yoon et al., 2024). While both methods employ negative prompts, their underlying mechanisms differ significantly. In SLD, the set of unsafe prompts, denoted as $c_{US}$, is designed to mitigate globally harmful image generation (Schramowski et al., 2023). In contrast, SAFREE focuses on more precise negative prompts specifically tailored to nudity-related content (Yoon et al., 2024). Beyond negative prompts, SAFREE further enhances safety by applying an orthogonal projection technique in Euclidean space to shift text embeddings away from predefined toxic regions. In the following, we provide an overview of the datasets used in our experiments.
@@
I2P The I2P dataset consists of prompts related to seven unsafe concepts: hate, harassment, violence, self-harm, sexual content, shocking content, and illegal activity (Schramowski et al., 2023). It contains a total of 4,703 prompts and was introduced in earlier stages of research, with subsequent studies primarily focusing on this dataset (Gong et al., 2024; Yoon et al., 2024). In this work, we utilize the I2P dataset as a source of reference data points rather than for additional training. The dataset was obtained from https://huggingface.co/datasets/AIML-TUDA/i2p
@@
Ring-A-Bell The Ring-A-Bell dataset was developed through a red-teaming approach that evaluates text-to-image diffusion models using black-box methods (Tsai et al., 2024). The original dataset Chia15/RingABell-Nudity contains 285 prompts; however, we use a curated subset of 79 prompts, following prior baselines (Gong et al., 2024; Yoon et al., 2024). This selection ensures a more equitable comparison of our method. The curated Ring-A-Bell dataset was obtained from either https://github.com/CharlesGong12/RECE or https://github.com/jaehong31/SAFREE.
@@
MMA-Diffusion MMA-Diffusion is another dataset generated via a red-teaming approach (Yang et al., 2024). Unlike other datasets, it consists of adversarial prompts designed to include potentially harmful contexts without explicit expressions. Similar to the Ring-A-Bell dataset, we use a curated set of 1,000 prompts, consistent with prior baselines (Gong et al., 2024; Yoon et al., 2024). The dataset was obtained from https://github.com/CharlesGong12/RECE or https://github.com/jaehong31/SAFREE.
@@
UnlearnDiff The UnlearnDiff dataset contains various harmful text prompts that can potentially generate NSFW images (Zhang et al., 2024b). Among its categories, we specifically focus on nudity-related prompts. The dataset includes a total of 116 nudity-related prompts, derived from an initial set of 143 prompts, from which 27 were excluded as they contained other NSFW categories such as self-harm and shocking content. This selection ensures that our numerical metrics remain unaffected by unrelated factors. The dataset was obtained from https://github.com/CharlesGong12/RECE or https://github.com/jaehong31/SAFREE.
@@
In Fig. 1, we demonstrate that SD-1.4 exhibits trainig dataset memorization, as it is capable of regenerating an indentical images using the text prompt, (’Living in the light with Ann Graham Lotz <|startoftext|> lad mans’). In this example, our method is applied with a bandwidth $\sigma=13.15$ and scaling factor of $0.69$. To construct a reference data for this case, we collected a total of 10 images from the internet. These are presented in Fig B.2.
Figure B.2: Reference images for Ann Graham Lotz case
@@
For unconditional generation, we utilize the FFHQ dataset to evaluate whether the proposed method effectively mitigates sexual bias, using our method. Although FFHQ datset does not include explicit label information, Table 3 illustrates that the generated images exibit a noticiable bias toward female images over male ones. To address this imbalance, we we use 1000 male images from CelebA-HQ [^5] https://www.kaggle.com/datasets/badasstechie/celebahq-resized-256x256 test dataset as reference data. While both FFHQ and CelebAHQ are designed to capture similar distribution, they are not completely aligned. This distinction provides an advantageous experimental setup, where we assess the controllability of image generation using reference images. For performance evaluation, we compute FID score using 1000 male images from the CelebA-HQ dataset. For classification tasks, we train a ResNet18 model, as implemented in the PyTorch framework [^6] https://pytorch.org/vision/stable/index.html using the CelebA-HQ training dataset.
@@
For conditional ImageNet (Russakovsky et al., 2015) experiments at $256\times 256$ resolution, we use a diffusion model trained on the full ImageNet-256 dataset guided by a classifier (Dhariwal & Nichol, 2021). The diffusion backbone uses a linear noise schedule across 1000 diffusion steps. We condition on class labels by scaling the classifier guidance at 5.0, creating a strong pull towards the desired class during the sampling process. Each experiment generates 50 samples per class across all 1000 ImageNet classes, producing 50,000 samples that are then evaluated with a pretrained ImageNet classifier for precision, recall, and classification accuracy measurements  (He et al., 2016b). Our metrics include (i) Precision: the fraction of generated samples that match the designated ImageNet label when conditioned on the class, (ii) Recall: aims to evaluate the diversity and coverage of the targeted class distribution, and (iii) Classification Accuracy: the rate at which generated images are correctly identified as their conditioned label among the 999 classes (excluding the negated target class, i.e, Chihuahua). The classification accuracy on the hold-out negated class is also calculated, to evaluate how well the respective method does not generate the negated target class. To avoid unintended Chihuahua generation, these metrics aim to make sure that samples do not drift toward distinct Chihuahua-like features when conditioning on other classes as well.
@@
For the experiments, we focus on the Chihuahua class to investigate how effectively our proposed safe denoiser can repel a target class while preserving generative quality for other classes. To compare our approach we implement three variants of the conditional diffusion process: a baseline classifier-guided diffusion model without repellency mechanisms, the Sparse Repellency (SR) (Kirchhof et al., 2024) technique applied to the classifier-guided diffusion model, and our safe denoiser technique applied to the same diffusion process. In this experiment, the safe denoiser technique is applied on the 200 to 800 timesteps of the diffusion process. A $\beta$ of $\beta=0.02$ was chosen as to control the strength of the repellency away from the Chihuahua target class. In the SR variant of the experiment, a repellency scale of 0.01 is combined with a large radius of 300 to push generated samples out of regions resembling the negated target class.
B.2 Additional Results

### v4

@@ 5.2 Class-Conditional and Unconditional Generation with Safe Denoiser
This subsection evaluates the performance of safe denoiser when applied in isolation. Specifically, to assess our safe denoiser in a simplified setting, we conduct experiments on two distinct tasks: a class-conditional model trained on ImageNet (russakovsky2015imagenet, ) for removing a specific class (e.g., Chihuahua); and on an unconditional model trained on FFHQ (karras2019style, ) to negate generating specific sex (e.g., female), see Appendix D for details of experiments. For the class removal, Table  4 presents precision, recall (kynkaanniemi2019improved, ), and Top-1* (classification accuracy of generated images conditioned by Chihuahua class) metrics for negating the Chihuahua class. As conventional text-based safety techniques are not directly applicable, we compare our method against Sparse Repellency (SR), as described in Sec. 4. Table  4 showcases that our method outperforms SR in terms of precision, recall, and Top-1*, indicating that ours avoid generating Chihuahua while being more diverse and precise than SR. In the FFHQ experiments, where we targeted the negation of female images, Table  6 indicates that while SR exhibits classification results that deceptively suggest successful negation, the FID scores and qualitative comparisons in Appendix E demonstrate that this apparent achievement comes at the cost of significantly degraded image quality. Indeed, our experiments show that our methodology consistently produces visually convincing samples, whereas SR frequently generates out-of-distribution images with artifacts.
6 Limitations and Conclusions
@@
We introduce the safe denoiser, a novel approach that modifies the sampling trajectories of DMs to adhere to theoretically safe distributions, thereby ensuring the generation of appropriate and authorized content. Experimental results demonstrate that the safe denoiser achieves state-of-the-art performance in concept erasing and class removal. Our method’s limitation is that it requires careful tuning of hyperparameters to balance image quality and safety. Appendix B discusses this further. Despite this, this work offers a robust and scalable solution for mitigating safety risks in generative AI, paving the way for responsible and ethical applications.
@@ Appendix A Proof
Appendix B Limitations and Broader Impacts
Limitations
We have addressed significant safety challenges in DMs, particularly concerning the generation of NSFW content and the inadvertent reproduction of sensitive data. We introduce the safe denoiser, a novel approach that modifies the sampling trajectories of DMs to adhere to theoretically safe distributions, thereby ensuring the generation of appropriate and authorized content.
However, this approach necessitates the introduction of an additional hyperparameter, $\beta_{t}$, as outlined in Theorem. 3.1. While we demonstrate that this parameter is theoretically derived and straightforward to implement, it may not be optimal for realistic scenarios due to its assumption of access to numerous data points sampled from an unsafe distribution. In practice, we present evidence in Figure 4(b) that this parameter influences the performance of the model.
Despite the challenges, we have developed a novel training-free method that effectively guides the sampling trajectories of DMs towards safe distributions. Ultimately, this work provides a robust and scalable solution for mitigating safety risks in generative AI, paving a way for their responsible and ethical applications.
Broader Impacts
Appendix C Experimental Details : Text-to-Image Generation
As outlined in the manuscript, we conduct the Text-to-Image experiment using SD-v1.4, following the same model as the baselines for generating images from text, as referenced in [2, 51, 12, 1]. To ensure consistency, we adopt the generation procedure described in each baseline. Preliminary observing the sensitivity of nudity-related content, we employ the DDPM scheduler [17]. For a fair comparison, we maintain the same number of inference steps, specifically $50$, aligning with the official implementations of both SLD and SAFREE, which also use 50 inference steps.
@@ Appendix C Experimental Details : Text-to-Image Generation
$K(x,x^{\prime})=\exp\left(-\frac{\lVert x-x^{\prime}\rVert^{2}}{2\sigma^{2}}\right)$ (C.1)
@@
For the bandwidth parameter $\sigma$, we set a value of 1.0 for SLD and 3.15 for SAFREE. Additionally, in case of SAFREE, we apply a scaling factor $\eta=0.33$, whereas for SLD, we use $\eta=0.03$ to regulate the strength of the repellency in Eq. (7). Empirically, we introduce a heuristic in which the proposed Safe Denoiser is applied within critical timesteps C=[780,,1000]C=[780,...,1000]. In the early stages of diffusion, denoising process primarily establishes global structures rather than intricate details, while the later stages focus on refining fine-grained features. Since our approach aims to prevent the generation of globally harmful images rather than enhancing image quality or detail, we apply the denoiser at these later timesteps.
For reference images, we provide a detailed explanation of how they are obtained. To ensure safe generation against nudity prompts, we utilize a total of 515 images sourced from the I2P dataset [2]. These images were generated using SD-v1.4. As mentioned in the manuscript, these reference images meet the criterion, where a nude class probability exceeds 0.6, as determined by Nudenet. Sample images are presented in Figure C.1. On the other hand, for the inappropriate probability task with the CoPro dataset, we attempt to use the total images from the I2P dataset. However, our computational resources allow us to use only 3,000 reference images. To select these 3,000 images, we randomly choose them out of the 4,703 images available in the I2P dataset. All images used in this task are also generated using SD-v1.4. Sample images are presented in Figure C.2. It’s important to note that all experiments conducted in this study use the same set of reference images across all baselines. This ensures a fair comparison.
Addtionally, the reference images we used in the task to generate safer images against nudity prompts are included as attachments in our supplementary materials. On the other hand, due to space constraints, we cannot include the attachment in the inappropriate probability task. We ensure that the reference images used in this task will be included in the public repository upon the acceptance of the paper.
Figure C.1: Reference images for safe generation against nudity prompts
Figure C.2: Reference images for inappropriate propability
@@
Next, we briefly introduce the baseline models used in our experiments. The first two approaches serve as comparisons for unlearning-based safe diffusion models [9, 12]. Specifically, we evaluate Erased Stable Diffusion (ESD) [9] as a representative method. More recently, reliably trained safe diffusion (RECE) models have demonstrated improved performance, particularly in reducing the attack success rate [12]. In addition to these unlearning-based approaches, we also include SLD and SAFREE as training-free safe diffusion models [2, 1]. While both methods employ negative prompts, their underlying mechanisms differ significantly. In SLD, the set of unsafe prompts, denoted as $c_{US}$, is designed to mitigate globally harmful image generation [2]. In contrast, SAFREE focuses on more precise negative prompts specifically tailored to nudity-related content [1]. Beyond negative prompts, SAFREE further enhances safety by applying an orthogonal projection technique in Euclidean space to shift text embeddings away from predefined toxic regions. In the following, we provide an overview of the datasets used in our experiments.
C.1 Inappropriate Prompt Datasets
@@
I2P The I2P dataset consists of prompts related to seven unsafe concepts: hate, harassment, violence, self-harm, sexual content, shocking content, and illegal activity [2]. It contains a total of 4,703 prompts and was introduced in earlier stages of research, with subsequent studies primarily focusing on this dataset [12, 1]. In this work, we utilize the I2P dataset as a source of reference data points rather than for additional training. The dataset was obtained from https://huggingface.co/datasets/AIML-TUDA/i2p
CoPro Compared to I2P [2], the CoPro dataset offers a more extensive dataset comprising a total of 226,104 prompts, each associated with 723 concepts that span both safe and unsafe scenarios. This expansion enhances the dataset’s suitability for rigorous evaluation [42]. Particularly, it also offers super-concept information, following the same framework of I2P [2]. All text prompts are categorized into {hate, harassment, violence, self-harm, sexual content, shocking content, and illegal activities}. This ensures that they align with the corresponding category information in the I2P dataset. To efficiently evaluate models, we randomly sample 10,000 prompts, ensuring a uniform distribution across all categories. We validated that the average inappropriate probability of SD-v1.4 in the randomly sampled dataset, presented in Table  3, closely matches the numerical information provided in [52]. In this work, we evaluate safe image generation performance across baselines on the CoPro dataset using reference data points from the I2P dataset. This dataset was obtained from https://github.com/rt219/LatentGuard/blob/main/dataset/CoPro_v1.0.json
C.2 Nudity in NSFW Prompt Datasets
@@
Ring-A-Bell The Ring-A-Bell dataset was developed through a red-teaming approach that evaluates text-to-image diffusion models using black-box methods [37]. The original dataset Chia15/RingABell-Nudity contains 285 prompts; however, we use a curated subset of 79 prompts, following prior baselines [12, 1]. This selection ensures a more equitable comparison of our method. The curated Ring-A-Bell dataset was obtained from either https://github.com/CharlesGong12/RECE or https://github.com/jaehong31/SAFREE.
@@
MMA-Diffusion MMA-Diffusion is another dataset generated via a red-teaming approach [26]. Unlike other datasets, it consists of adversarial prompts designed to include potentially harmful contexts without explicit expressions. Similar to the Ring-A-Bell dataset, we use a curated set of 1,000 prompts, consistent with prior baselines [12, 1]. The dataset was obtained from https://github.com/CharlesGong12/RECE or https://github.com/jaehong31/SAFREE.
@@
UnlearnDiff The UnlearnDiff dataset contains various harmful text prompts that can potentially generate NSFW images [36]. Among its categories, we specifically focus on nudity-related prompts. The dataset includes a total of 116 nudity-related prompts, derived from an initial set of 143 prompts, from which 27 were excluded as they contained other NSFW categories such as self-harm and shocking content. This selection ensures that our numerical metrics remain unaffected by unrelated factors. The dataset was obtained from https://github.com/CharlesGong12/RECE or https://github.com/jaehong31/SAFREE.
C.3 Ann Graham Lotz for Data Memorization
@@
In Figure 1(b), we demonstrate that SD-v1.4 exhibits trainig dataset memorization, as it is capable of regenerating an indentical images using the text prompt, (’Living in the light with Ann Graham Lotz <|startoftext|> lad mans’). In this example, our method is applied with a bandwidth $\sigma=13.15$ and scaling factor of $0.69$. To construct a reference data for this case, we collected a total of 10 images from the internet. These are presented in Figure C.3
Figure C.3: Reference images for Ann Graham Lotz case
Appendix D Experimental Details : Class-Conditional and Unconditional Generation
In this section, we use our safe denoiser in the DMs without text inputs. Specifically, we employ experiments on FFHQ [49] and ImageNet [48] in the $256\times 256$ resolution. We utilize the pretrained diffusion models from FFHQ [53] [^6] https://github.com/DPS2022/diffusion-posterior-sampling and ImageNet [7] [^7] https://github.com/openai/guided-diffusion. For the experiments, we use a DDPM solver [54] with 100 steps.
@@ Appendix D Experimental Details : Class-Conditional and Unconditional Generation
For unconditional generation, we utilize the FFHQ dataset to evaluate whether the proposed method effectively mitigates sexual bias, using our method. Although FFHQ datset does not include explicit label information, Table  6 illustrates that the generated images exibit a noticiable bias toward female images over male ones. In this experiment, we select 1K female images from CelebA-HQ [55] [^8] https://www.kaggle.com/datasets/badasstechie/celebahq-resized-256x256 validation split to serve as unseen negative data, thereby establishing the negative dataset $\{\mathbf{x}^{(1)},...,\mathbf{x}^{(1000)}\}$. Then, we employ our safe denoiser to generate 1K images. While both FFHQ and CelebA-HQ are designed to capture similar distribution, they are not completely aligned. This distinction provides an advantageous experimental setup, where we assess the controllability of image generation using reference images. For performance evaluation, we compute FID [43] score using 1,000 male images from the CelebA-HQ dataset. For classification performance, we train a ResNet18 model, as implemented in the PyTorch framework  [^9] https://pytorch.org/vision/stable/index.html using the training dataset in the CelebA-HQ. In this experiment, we chose $\sigma=1.0$ and $\eta=0.05$, and employ Safe denoiser across the entire denoising timesteps.
@@
For conditional ImageNet [48] experiments at $256\times 256$ resolution, we use a diffusion model trained on the full ImageNet-256 dataset guided by a classifier [7]. The diffusion backbone follows a linear noise schedule and is constructed with 1,000 diffusion time-steps. We condition on class labels by scaling the classifier guidance at 5.0, creating a strong pull towards the desired class during the sampling process. Each experiment generates 50 samples per class across all 1000 ImageNet classes, producing 50,000 samples that are then evaluated with a pretrained ImageNet classifier for precision, recall, and classification accuracy measurements [56]. Our metrics include (i) Precision: the fraction of generated samples that match the designated ImageNet label when conditioned on the class, (ii) Recall: aims to evaluate the diversity and coverage of the targeted class distribution, and (iii) Classification Accuracy: the rate at which generated images are correctly identified as their conditioned label among the 999 classes (excluding the negated target class, i.e, Chihuahua). The classification accuracy on the hold-out negated class is also calculated, to evaluate how well the respective method does not generate the negated target class. As illustrated in Table  4, we focus on the Chihuahua class to investigate how effectively our proposed safe denoiser can repel a target class while preserving generative quality for other classes in this experiment. To avoid unintended Chihuahua generation, aforementioned metrics aim to make sure that samples do not drift toward distinct Chihuahua-like features. For instance, when we generate an image based on a reference dataset sampled from a Chihuahua, the resulting sample may resemble a Golden Retriever, but it won’t resemble a Chihuahua.
@@
To compare our approach, we implement three variants of the conditional diffusion process: vanilla classifier-guided diffusion model without repellency mechanisms, the Sparse Repellency (SR) [38] technique applied to the classifier-guided diffusion model, and our safe denoiser technique applied to the same diffusion process. For the reference dataset, we select the validation set of Chihuahua class as the negative images. In this experiment, the safe denoiser technique is applied on the 200 to 800 timesteps of the diffusion process. $\eta=0.02$ was chosen as to control the strength of the repellency away from the Chihuahua target class. In the SR variant of the experiment, a repellency scale of 0.01 is combined with a large radius of 300 to push generated samples out of regions resembling the negated target class.
Appendix E Additional Experimental Results
In this section, we share extra experimental results. Both numeric and visual results are included, which are not presented in the main text. These results highlight the empirical gains in terms of safe generation and the preservation of the global context of the generated samples simultaneously. Specifically, this ensures that the samples remain faithful to their original meanings while enabling us to negate specific concepts we intended.
E.1 Quantiative Results
ImageNet Case for Negating Chihuahua Class
In ImageNet, we focus on negating a specific Chihuahua class during generation. We select the validation set of Chihuahua class as the negative images. We generate 50 samples per class and classify samples from 999 classes by a classifier [7] and report the accuracy by Top-1. Also, we measure the Top-1 accuracy of 50 samples from Chihuahua class, reporting it by Top-1* in Table  E.1. From the result, we note that our method excels generating other 999 classes, while SR cannot generate images from those 999 classes. To evaluate the overall quality, Table  E.1 further report the precision (sample accuracy) and recall (sample diversity) [50] over 50K samples, indicating that our method is better than SR in negating a specific class.
Table E.1: Experiments on ImageNet for the specific class (Chihuahua) negation task. Top-1 is the classification accuracy of the generated samples on 999 classes, and Top-1* indicates the accuarcy on the specific class.
Method Prec $\uparrow$ Rec $\uparrow$ Top-1 $\uparrow$ Top-1 $\downarrow$
Baseline (B) 0.72 0.63 0.76 0.68
B + SR 0.59 0.54 0.01 0.0
B + Ours 0.62 0.58 0.14 0.0
Aesthetic Scores for Long Text Prompts
We identified that our method, which incorporates negative prompts, effectively reduces the risk of generating unsafe data and maintains alignment with the given text prompts. However, the text prompts in these cases span various lengths. Therefore, it is necessary to quantify whether our method excessively applies to remove unsafe contents, leading to unfavorable images in extreme cases. We sample the most complex cases from the I2P datasets and compare the generated images across baselines.
Table E.2: Aesthetic scroes for long text prompts in the I2P dataset.
Method LAION-aesthetic V2 $\uparrow$
SD-v1.4 $5.97\pm 0.534$
SAFREE $6.03\pm 0.540$
SAFREE+Ours $5.94\pm 0.529$
To test how our method works when long and complex prompts are given, we use LAION-aesthetic V2 score  [^10] https://github.com/christophschuhmann/improved-aesthetic-predictor as a metric and use top 10% longest prompts (475 prompts, avg. word_count=54) selected from the I2P dataset. We choose this score as it is known to be correlated with human perception of quality of images (higher the better). As shown Table  E.2, our method maintains aesthetic quality comparable to baselines, even with complex prompts. We prove that the proposed method does not struggle to create appropriate images even when asked with long text prompts.
E.2 Qualitative Results